The solution to a problem is represented by the following equation:

(9.0×109N⋅m2/C2)(+20.0×10−9C)(+50.0×10−9C)/(5cm)=12(10g)v2x+(9.0×109N⋅m2/C2)(+20.0×10−9C)(+50.0×10−9C)/(20cm)


Describe a situation that the equation might represent and formulate a problem for which it is a solution (there are multiple possibilities).
Answer #1

Using conservation of energy, we have

P.Einitial + K.Einitial = K.Efinal + P.Efinal

(ke q1 q2 / ri) + 0 = (1/2) m v2 + (ke q1 q2 / rf)

[(9 x 109 Nm2/C2) (20 x 10-9 C) (50 x 10-9 C)] / (0.05 m) = (1/2) (0.01 kg) v2 + (9 x 109 Nm2/C2) (20 x 10-9 C) (50 x 10-9 C)] / (0.2 m)

[(9 x 10-6 Nm2) / (0.05 m)] = (0.005 kg) v2 + [(9 x 10-6 Nm2) / (0.2 m)]

(1.80 x 10-4 J) = (0.005 kg) v2 + (4.50 x 10-5 J)

[(1.80 x 10-4 J) - (4.50 x 10-5 J)] = (0.005 kg) v2

[(1.35 x 10-4 J) / (0.005 kg)] = v2

v = \sqrt{}0.027 m2/s2

v = 0.164 m/s